OpenStudy (anonymous):

A. y + 2 =3/4 (x - 2) B. y - 1 =3/4 (x - 2) C. y + 2 =4/3 (x + 2) D. y -1 =3/4 (x + 2) E. y - (-2) =3/4 (x - (-2)) F. y + 2 =3/4 (x + 2) Which of the point-slope equations below are correct for the line that passes through points (-2, -2) and (2, 1)?

OpenStudy (anonymous):

f

OpenStudy (anonymous):

only F?

OpenStudy (anonymous):

It seems that E is the same as F

OpenStudy (anonymous):

(y-(-2))/(x-(-2))=(1-(-2))/(2-(-2)) (y+2)/(x+2)=3/4 Hence Both E and F

OpenStudy (anonymous):

Elias and rs are correct. e and f are the same and they are the only 2 answers

OpenStudy (anonymous):

Agree with @thomasj